OpenStudy (anonymous):

solve the equation 5(2x)+1=3(5x-4) x=???

OpenStudy (anonymous):

5(2x)+1=3(5x-4) 10x+1=3(5x)-3(4) 10x+1=15x-12 10x+1-10x=15x-12-10x 1=5x-12 5x-12=1 5x-12+12=1+12 5x=13 5x/5=13/5 x=13/5